OpenStudy (anonymous):

Someone please help me! Im stuck! Medal+Fan A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (vlery):

couldn't you just solve for either x or Y and then put that answer into the bottom equation?

OpenStudy (anonymous):

x=7y/3 + 16/3

OpenStudy (anonymous):

y=3x/7 - 16/7

OpenStudy (vlery):

-7y/3-16/3

OpenStudy (anonymous):

Would that be it?? Im confused....

OpenStudy (anonymous):

?

OpenStudy (vlery):

no now you plug that into the next equation where ever there is an (x)

OpenStudy (anonymous):

How would I do that...? Sorry :(

OpenStudy (vlery):

-9x + 5y = 16 but now put x=-7y/3-16/3 into where ever x is

OpenStudy (anonymous):

-9-7y/3-16/3y=17 ?

OpenStudy (vlery):

-9(-7y/3-16/3)+5y=16

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Then what?

OpenStudy (vlery):

distribute -9 into -7y/3-16/3

OpenStudy (anonymous):

To where, what number?

OpenStudy (vlery):